OpenStudy (anonymous):

Help with Multiplying Rational Expressions: x^2-x-6/3x-6 • x-2/x^2+x-2

OpenStudy (luigi0210):

start by factoring

OpenStudy (luigi0210):

\[\frac{ x^2-x-6 }{ 3x-6 }*\frac{ x-2 }{ x^2+x-2 }\]

OpenStudy (luigi0210):

\[\frac{ (x-3)(x+2) }{ 3(x-2) }*\frac{ x-2 }{ (x-1)(x+2) }\] Can you handle it from here?

OpenStudy (luigi0210):

Well first cancel like terms, then just multiply the remaining terms

OpenStudy (luigi0210):

where did the 4 come from?